# Boundary transfer matrices arising from quantum symmetric pairs

### v1

@@ 5.6. QSP weight modules
The tensor K-matrix associated to $\Upsilon$ does not automatically act on any $U_{q}(\mathfrak{k})$-module. We need to specify a suitable category of weight $U_{q}(\mathfrak{k})$-modules , see [AV24] and cf. [BW18a, Wa24]. Set
@@
with the canonical projection $[\cdot]_{\phi}:\mathsf{P}\twoheadrightarrow\mathsf{P}_{\phi}$. There exists a pairing $\mathsf{P}_{\phi}\times\mathsf{Q}^{\vee}_{\sf ext}\to\tfrac{1}{2}{\mathbb{Z}}$ defined by $\zeta(h)=\lambda(h+\phi(h))/2$ if $\zeta=[\lambda]_{\phi}\in\mathsf{P}_{\phi}$. The restriction to $\mathsf{P}_{\phi}\times(\mathsf{Q}^{\vee}_{\sf ext})^{\phi}$ is nondegenerate and takes values in ${\mathbb{Z}}$. Note that $U_{q}({\mathfrak{h}}^{\phi})={\mathbb{F}}\langle\{k_{h}\}_{h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\phi}}\rangle$.
@@
We denote the full subcategory of QSP weight modules by $\mathcal{W}_{\phi}\subseteq{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ and the corresponding completion of $U_{q}(\mathfrak{k})$ by $\operatorname{End}({\mathsf{f}}_{\phi})=\operatorname{End}({\mathsf{f}}_{\mathcal{W}_{\phi}})$. [^15] Note that $\mathcal{W}_{\phi}={\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ if and only if $\phi|_{{\mathfrak{h}}}=\omega|_{{\mathfrak{h}}}$. The category $\mathcal{W}_{\phi}$ has many desirable properties, see [AV24, Prop. 5.2.3]:
@@ 5.7. Tensor K-matrices for quantum symmetric pairs
Let ${{\mathbb{K}}}$ be the tensor K-matrix corresponding to $K=\Upsilon$.
Theorem 5.7.1 ([AV24]).
@@
The proof relies on the maximality property (2.19), see [AV24, Thm. 5.6.1], which holds as a consequence of Proposition 5.2.3 (2), a study of the classical limit of $\Upsilon$, and the Iwasawa decomposition (5.5). Note that, for QSPs of finite type and finite-dimensional $U_{q}(\mathfrak{k})$-modules, the support property is proved in [Kol20].
@@ 5.8. The boundary transfer matrix for the dual K-matrix
It is enough to apply the proof of Theorem 2.5.2, although the algebraic manipulations needs to be understood in $\operatorname{End}({\mathsf{f}})$. To see that $\overline{\mathbb{K}}_{\bullet,V}\in U_{q}(\mathfrak{k})\otimes\operatorname{End}(V)$ (and hence $\mathscr{T}_{\phi}^{(V)}\in U_{q}(\mathfrak{k})$), it is convenient to factorize ${\overline{{\mathbb{K}}}}$. Recall the factorization $R=\kappa\cdot\Xi$ from Section 4.4. For a ${\mathbb{Z}}$-linear function $f:\mathsf{P}\to\mathsf{P}$, consider $\kappa^{f}=(\kappa^{f}_{V,W})_{V,W\in\mathcal{W}}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{2})$ with $\kappa^{f}_{V,W}$ acting on $V_{\lambda}\otimes W_{\mu}$ as multiplication by $q^{(f(\lambda),\mu)}$ (so $\kappa=\kappa^{\operatorname{id}}$). Then
@@
Since $V\in{\mathcal{O}}$, $\Upsilon_{V}\in\operatorname{End}(V)$ and $\Xi_{\bullet,V}\in U_{q}({\mathfrak{g}})\otimes\operatorname{End}(V)$. Moreover, since $V$ is finite-dimensional, one shows that $(\Xi_{V^{\psi},\bullet})_{21}\in U_{q}({\mathfrak{g}})\otimes\operatorname{End}(V)$, see [AV24, Thm. 5.6.1]. Therefore, ${\overline{{\mathbb{K}}}}_{\bullet,V}\in U_{q}(\mathfrak{k})\otimes\operatorname{End}(V)$. ∎
@@
The image of $\mathscr{T}_{\phi}$ lies in $U_{q}({\mathfrak{g}}_{X})U_{q}({\mathfrak{h}}^{\phi})$.
@@
$\mathscr{T}_{\phi}^{(V)}=\mathop{\sf Tr}_{V}(1\otimes D_{V})\cdot\kappa^{\phi}_{\bullet,V}\cdot\kappa_{\bullet,V}=\mathop{\sf Tr}_{V}(1\otimes D_{V})\cdot\kappa^{{\operatorname{id}}+\phi}_{\bullet,V}.$ (5.18)
@@ 5.9. The boundary transfer matrix for the Drinfeld element
We now assume that $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ is a QSP. In this case, the standard cylindrical structure is gauge equivalent to $(\varphi,R_{21}^{-1},K^{\sf BK})$ where $\varphi\colon U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$ is a quasitriangular Hopf algebra automorphism, and $K^{\sf BK}\in\operatorname{End}({\mathsf{f}})$ is the basic K-matrix considered in [BK19, Sec. 7.3].
@@
    From this point of view, it is convenient to consider the semistandard cylindrical structure on $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$, which has the form $(\omega\circ\tau,1\otimes 1,K)$ where $K\in\operatorname{End}({\mathsf{f}})$ is the semistandard K-matrix, see [AV22a]. Thus, we obtain a boundary transfer matrix map by setting $K^{\prime}$ equal to the semistandard K-matrix of an arbitrary QSP subalgebra whose underlying diagram automorphism is $\tau$. $\triangledown$
@@ 6. Boundary transfer matrices for quantum affine symmetric pairs
Henceforth, assume that our QSP $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ is of untwisted affine type. We are actually interested in finite-dimensional modules of untwisted quantum loop algebras, viewing the ring homomorphisms $\mathscr{T}_{\phi}$ as refined representation-theoretic tools, in the spirit of [FR99]. We will define spectral analogues of $\mathscr{T}^{(V)}_{\phi}$. The simplification explained in Proposition 5.8.2 does not occur since the relevant matrices are not all triangular in the weight bases of these modules.
@@ 6.1. Affine Lie algebras [Kac90]
Let $c\in{\mathfrak{h}}^{\prime}$ be the canonical central element and $h^{\vee}=\rho(c)$ the dual Coxeter number. We have the linear decomposition ${\mathfrak{h}}^{\prime}\cong\mathring{{\mathfrak{h}}}\oplus{\mathbb{C}}c$ with associated projection ¯$\overline{\phantom{\imath}\cdot\phantom{\imath}}:{\mathfrak{h}}^{\prime}\to\mathring{{\mathfrak{h}}}$. Denote the monoidal category of finite-dimensional type-${\bf 1}$ $U_{q}({\mathfrak{g}}^{\prime})$-modules by ${\mathcal{C}}$. For all $V\in{\mathcal{C}}$ we have the decomposition
@@ 6.3. Crossing symmetry
$R^{s}_{V,{}^{**}W}(z)=(((R^{s}_{V,W}(z)^{-1})^{\mathsf{t}_{2}})^{-1})^{\mathsf{t}_{2}}.$ (6.6)
The specific grading shift we need is the $\tau$-minimal grading shift and is defined in terms of the diagram automorphism $\tau$ associated to $U_{q}(\mathfrak{k})$. Namely, we define $s\in\operatorname{Hom}_{\sf grp}(\mathsf{Q},{\mathbb{Z}})$ by
@@
Clearly, $s=s_{\sf hom}$ if and only if $\tau(0)=0$. We will abbreviate
@@
On the other hand, let $\Sigma^{\sf hom}=\Sigma^{s_{\sf hom}}$ be the homogeneous grading shift, so $s_{\sf hom}(\alpha_{i})=\delta_{i0}$, and consider the corresponding spectral R-matrix $R^{\sf hom}(z)=({\operatorname{id}}\otimes\Sigma^{\sf hom}_{z})(R)$. It is convenient to have a conversion rule for these two grading shifts and the corresponding spectral R-matrices. Set
@@
$R_{V,W}(z)=\operatorname{Ad}(\operatorname{Id}_{V}\otimes z^{s-f\,{s_{\sf hom}}})(R^{\sf hom}_{V,W}(z^{f}))=\operatorname{Ad}(z^{f\,s_{\sf hom}-s}\otimes\operatorname{Id}_{W})(R^{\sf hom}_{V,W}(z^{f})).$ (6.10)
Denote the sum of fundamental $\mathring{{\mathfrak{g}}}$-weights by $\mathring{\rho}$ and extend it to an element of $({\mathfrak{h}}^{\prime})^{*}$ by setting $\mathring{\rho}(c)=0$. Define a linear map $\mathring{D}_{V}$ on $V\in{\mathcal{C}}$ via
$\mathring{D}_{V}(v)=q^{-2(\mathring{\rho},\lambda)+2(h^{\vee}s_{\sf hom}-h^{\vee}_{\phi}s)(\lambda)}v,\qquad v\in V_{\lambda},\qquad\lambda\in\mathring{\mathsf{P}}$ (6.11)
and collect them into the tuple $\mathring{D}=(\mathring{D}_{V})_{V\in{\mathcal{C}}}\in\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$. Then, for any root vector $x\in U_{q}({\mathfrak{g}})$ of weight $\alpha\in\mathsf{Q}$ we have
$\operatorname{Ad}(\mathring{D})(x)=q^{-2(\mathring{\rho},\alpha)+2(h^{\vee}s_{\sf hom}-h^{\vee}_{\phi}s)(\alpha)}x.$ (6.12)
@@
Set $p=q^{-h_{\tau}^{\vee}}$. The action of $R(z)$ on tensor products in ${\mathcal{C}}$ satisfies
@@
We first note that $\rho=h^{\vee}\Lambda_{0}+\mathring{\rho}\in({\mathfrak{h}}^{\prime})^{*}$, where $\Lambda_{0}\in({\mathfrak{h}}^{\prime})^{*}$ is defined by $\Lambda_{0}(h_{i})=\delta_{i,0}$ for all $i\in I$, see e.g. [Kac90, Ex. 7.16]. By a straightforward check on generators using (6.12), one obtains $\operatorname{Ad}(q^{2(\rho-\mathring{\rho})})=\Sigma^{\sf hom}_{q^{2h^{\vee}}}$. This in turn implies the following statements for the homogeneous grading-shifted R-matrix:
$({\operatorname{id}}_{V}\otimes\operatorname{Ad}(q^{2(\rho-\mathring{\rho})})(R^{\sf hom}_{V,W}(q^{-2h^{\vee}}z^{f}))=R^{\sf hom}_{V,W}(z^{f}).$ (6.14)
Next, we use (6.10) to obtain analogous statements in the general case:
$(\operatorname{Id}_{V}\otimes\operatorname{Ad}(q^{2(\mathring{\rho}-\rho)}q^{2(h^{\vee}_{\phi}s-h^{\vee}s_{\sf hom})})(R_{V,W}(p^{2}z))$ $=R_{V,W}(z).$ (6.15)
These are equivalent to the desired identities since $\operatorname{Ad}(\mathring{D})=S^{2}$. ∎

### v2

@@ 5.6. QSP weight modules
The tensor K-matrix associated to $\Upsilon$ does not act on any tensor product of a $U_{q}(\mathfrak{k})$-module and a $U_{q}({\mathfrak{g}})$-module. We need to specify a suitable category of weight $U_{q}(\mathfrak{k})$-modules, see [AV25b] and cf. [BW18b, Wa24]. Set
@@
with the canonical projection $\cdot]_{\phi}:\mathsf{P}\twoheadrightarrow\mathsf{P}_{\phi}$. There exists a pairing $\mathsf{P}_{\phi}\times\mathsf{Q}^{\vee}_{\sf ext}\to\tfrac{1}{2}{\mathbb{Z}}$ defined by $\zeta(h)=\lambda(h+\phi(h))/2$ if $\zeta=[\lambda]_{\phi}\in\mathsf{P}_{\phi}$. The restriction to $\mathsf{P}_{\phi}\times(\mathsf{Q}^{\vee}_{\sf ext})^{\phi}$ is nondegenerate and takes values in $\mathbb{Z}$}. Note that $U_{q}({\mathfrak{h}}^{\phi})={\mathbb{F}}\langle\{k_{h}\}_{h\in(\mathsf{Q}^{\vee}_{\sf ext})^{\phi}}\rangle$.
@@
We denote the full subcategory of QSP weight modules by $\mathcal{W}_{\phi}\subseteq{\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ and the corresponding completion of $U_{q}(\mathfrak{k})$ by $\operatorname{End}({\mathsf{f}}_{\phi})=\operatorname{End}({\mathsf{f}}_{\mathcal{W}_{\phi}})$. [^15] Note that $\mathcal{W}_{\phi}={\operatorname{Mod}}(U_{q}(\mathfrak{k}))$ if and only if $\phi|_{{\mathfrak{h}}}=\omega|_{{\mathfrak{h}}}$. The category $\mathcal{W}_{\phi}$ has many desirable properties, see [AV25b, Prop. 5.2.3]:
@@ 5.7. Tensor K-matrices for quantum symmetric pairs
Let $\mathbb{K}$}} be the tensor K-matrix corresponding to $K=\Upsilon$, i.e., ${{\mathbb{K}}}\coloneqq(R^{\psi})_{21}\cdot\Upsilon_{2}\cdot R$.
Theorem 5.7.1 ([AV25b]).
@@
The proof relies on the maximality property (2.19), see [AV25b, Thm. 5.6.1], which holds as a consequence of Proposition 5.2.3 (2), a study of the classical limit of $\Upsilon$, and the Iwasawa decomposition (5.5). For QSPs of finite type and finite-dimensional $U_{q}(\mathfrak{k})$-modules, the support property is proved in [Kol20].
Remark 5.7.2.
In contrast with Section 5.5, tensor K-matrices of finite types are not known to decompose according to the restricted root system. For QSPs of split and quasi-split affine types, new loop generators have been introduced in [LW21, Zh22, LWZ23, LWZ24]. No loop-triangular decomposition is known for their tensor K-matrices. However, for $U_{q}(\widehat{\mathfrak{sl}}_{N})$, strong evidence is given by the K-matrix presentations of the QSPs in [MRS03, CGM14]. On a related point, the authors of [LBG25a] provide a twisted LDU decomposition of an algebra-valued matrix which should correspond to $\mathbb{K}_{\bullet,V}(z)$ in the current work, in the case where $U_{q}(\mathfrak{k})\subset U_{q}(\widehat{\mathfrak{sl}}_{2})$ is the q-Onsager algebra and $V$ is the two-dimensional irreducible $U_{q}(\widehat{\mathfrak{sl}}_{2})$-module. $\triangledown$
@@ 5.8. The boundary transfer matrix for the dual K-matrix
It is enough to apply the proof of Theorem 2.5.2, although the algebraic manipulations need to be understood in $\operatorname{End}({\mathsf{f}})$. To see that $\overline{\mathbb{K}}_{\bullet,V}\in U_{q}(\mathfrak{k})\otimes\operatorname{End}(V)$ (and hence $\mathscr{T}_{\phi}^{(V)}\in U_{q}(\mathfrak{k})$), it is convenient to factorize $\overline{\mathbb{K}}$. Recall the factorization $R=\kappa\cdot\Xi$ from Section 4.4. For a $\mathbb{Z}$}-linear function $f:\mathsf{P}\to\mathsf{P}$, consider $\kappa^{f}=(\kappa^{f}_{V,W})_{V,W\in\mathcal{W}}\in\operatorname{End}({\mathsf{f}}_{\mathcal{W}}^{2})$ with $\kappa^{f}_{V,W}$ acting on $V_{\lambda}\otimes W_{\mu}$ as multiplication by $q^{(f(\lambda),\mu)}$ (so $\kappa=\kappa^{\operatorname{id}}$). Then
@@
Since $V\in{\mathcal{O}}$, $\Upsilon_{V}\in\operatorname{End}(V)$ and $\Xi_{\bullet,V}\in U_{q}({\mathfrak{g}})\otimes\operatorname{End}(V)$. Moreover, since $V$ is finite-dimensional, one shows that $(\Xi_{V^{\psi},\bullet})_{21}\in U_{q}({\mathfrak{g}})\otimes\operatorname{End}(V)$, see [AV25b, Thm. 5.6.1]. Therefore, $\overline{\mathbb{K}}_{\bullet,V}\in U_{q}(\mathfrak{k})\otimes\operatorname{End}(V)$. ∎
@@
The image of $\mathscr{T}_{\phi}$ lies in $U_{q}({\mathfrak{g}}_{X})U_{q}({\mathfrak{h}}^{\phi})^{\mathsf{P}}$, where $U_{q}({\mathfrak{h}}^{\phi})^{\mathsf{P}}$ is the weight lattice refinement of $U_{q}({\mathfrak{h}}^{\phi})$.
@@
$\mathscr{T}_{\phi}^{(V)}=\mathop{\sf Tr}_{V}((1\otimes D_{V})\cdot\kappa^{{\operatorname{id}}+\phi}_{\bullet,V})=\sum_{\mu\in\mathsf{P}}\dim(V_{\mu})k_{\mu+\phi(\mu)-2\rho}$ (5.17)
where $k_{\lambda}|_{V_{\mu}}=q^{(\lambda,\mu)}$. ∎
@@ 5.9. The boundary transfer matrix for the Drinfeld element
We now assume that $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ is a QSP of finite type. In this case, the standard cylindrical structure is gauge equivalent to $(\varphi,R_{21}^{-1},K^{\sf BK})$ where $\varphi\colon U_{q}({\mathfrak{g}})\to U_{q}({\mathfrak{g}})$ is a quasitriangular Hopf algebra automorphism, and $K^{\sf BK}\in\operatorname{End}({\mathsf{f}})$ is the basic K-matrix considered in [BK19, Sec. 7.3].
@@
    From this point of view, it is convenient to consider the semistandard cylindrical structure on $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$, which has the form $(\omega\circ\tau,1\otimes 1,K)$ where $K\in\operatorname{End}({\mathsf{f}})$ is the semistandard K-matrix, see [AV22]. Thus, we obtain a boundary transfer matrix map by setting $K^{\prime}$ equal to the semistandard K-matrix of an arbitrary QSP subalgebra whose underlying diagram automorphism is $\tau$. $\triangledown$
@@ 6. Boundary transfer matrices for quantum affine symmetric pairs
Henceforth, assume that our QSP $(U_{q}({\mathfrak{g}}),U_{q}(\mathfrak{k}))$ is of untwisted affine type. We are actually interested in finite-dimensional modules of untwisted quantum loop algebras, viewing the ring homomorphisms $\mathscr{T}_{\phi}$ as refined representation-theoretic tools, in the spirit of [FR99]. We will define spectral analogues of $\mathscr{T}^{(V)}_{\phi}$. The simplification explained in Lemma 5.8.2 does not occur since the relevant matrices are not all triangular in the weight bases of these modules.
@@ 6.1. Affine Lie algebras [Kac90]
Let $c\in{\mathfrak{h}}^{\prime}$ be the canonical central element and $h^{\vee}=\rho(c)$ the dual Coxeter number. We have the linear decomposition ${\mathfrak{h}}^{\prime}\cong\mathring{{\mathfrak{h}}}\oplus{\mathbb{C}}c$ with associated projection $\overline{\phantom{\imath}\hskip-4.0pt\cdot\hskip-4.0pt\phantom{\imath}}:{\mathfrak{h}}^{\prime}\to\mathring{{\mathfrak{h}}}$. Denote the monoidal category of finite-dimensional type-$\bf 1$} $U_{q}({\mathfrak{g}}^{\prime})$-modules by $\mathcal{C}$}. For all $V\in{\mathcal{C}}$ we have the decomposition
@@ 6.3. Crossing symmetry
$\hskip-4.0ptR^{s}_{V,{}^{**}W}(z)=(((R^{s}_{V,W}(z)^{-1})^{\mathsf{t}_{2}})^{-1})^{\mathsf{t}_{2}}.$ (6.6)
The specific grading shift we need is the $\tau$-minimal grading shift and is defined in terms of $\tau$, the diagram automorphism factor of the Lie algebra automorphism $\phi$ used to define $U_{q}(\mathfrak{k})$. Namely, we define $s\in\operatorname{Hom}_{\sf grp}(\mathsf{Q},{\mathbb{Z}})$ by
@@
We will abbreviate
@@
On the other hand, let $\Sigma^{\sf hom}=\Sigma^{s_{\sf hom}}$ be the homogeneous grading shift, so $s_{\sf hom}(\alpha_{i})=\delta_{i0}$ (clearly, $s=s_{\sf hom}$ if and only if $\tau(0)=0$), and consider the corresponding spectral R-matrix $R^{\sf hom}(z)=({\operatorname{id}}\otimes\Sigma^{\sf hom}_{z})(R)$. It is convenient to have a conversion rule for these two grading shifts and the corresponding spectral R-matrices. Set
@@
$R_{V,W}(z)$ $=\operatorname{Ad}(\operatorname{Id}_{V}\otimes z^{s-f\,{s_{\sf hom}}})(R^{\sf hom}_{V,W}(z^{f})).$ (6.10)
Denote the sum of fundamental $\mathring{{\mathfrak{g}}}$-weights by $\mathring{\rho}$ and extend it to an element of $({\mathfrak{h}}^{\prime})^{*}$ by setting $\mathring{\rho}(c)=0$. Define linear maps $\mathring{D}_{V}:V\to V$ ($V\in{\mathcal{C}}$) via
$\mathring{D}_{V}(v)=q^{-2(\mathring{\rho},\lambda)+2(h^{\vee}_{\phi}s-h^{\vee}s_{\sf hom})(\lambda)}v,\qquad v\in V_{\lambda},\qquad\lambda\in\mathring{\mathsf{P}}$ (6.11)
and collect them into the tuple $\mathring{D}=(\mathring{D}_{V})_{V\in{\mathcal{C}}}\in\operatorname{End}({\mathsf{f}}_{{\mathcal{C}}})$.
@@
Set $p=q^{h_{\phi}^{\vee}}$. The action of $R(z)$ on tensor products in $\mathcal{C}$} satisfies
@@
We first review the proof in the case of the homogeneous grading. By combining (6.6) and $S^{-2}=\operatorname{Ad}(q^{2\rho})$ we obtain
$(((R^{\sf hom}_{V,W}(z)^{-1})^{\mathsf{t}_{2}})^{-1})^{\mathsf{t}_{2}}=\big({\operatorname{id}}\otimes\operatorname{Ad}(q^{2\rho})\big)\big(R^{\sf hom}_{V,W}(z)\big).$ (6.13)
Next, we recall that $\rho=\mathring{\rho}+h^{\vee}\Lambda_{0}$, where $\Lambda_{0}\in({\mathfrak{h}}^{\prime})^{*}$ is defined by $\Lambda_{0}(h_{i})=\delta_{i,0}$ for all $i\in I$, see e.g. [Kac90, Ex. 7.16]. By a straightforward check on generators, one obtains $\operatorname{Ad}(q^{2\rho})=\operatorname{Ad}(q^{2\mathring{\rho}})\circ\Sigma^{\sf hom}_{q^{2h^{\vee}}}$. Hence
$(((R^{\sf hom}_{V,W}(z)^{-1})^{\mathsf{t}_{2}})^{-1})^{\mathsf{t}_{2}}=\big({\operatorname{id}}\otimes\operatorname{Ad}(q^{2\mathring{\rho}})\big)\big(R^{\sf hom}_{V,W}(q^{2h^{\vee}}z)\big),$ (6.14)
which is the desired identity for the homogeneous grading. The identity for the $\tau$-minimal grading readily follows by replacing $z$ by $z^{f}$, conjugating by $\operatorname{Id}_{V}\otimes z^{s-f\,s_{\sf hom}}$, and applying the conversion rules (6.10). ∎